### v1

# Neural Risk Limiting Dispatch in Power Networks: Formulation and Generalization Guarantees

Abstract
Risk limiting dispatch (RLD) has been proposed as an approach that effectively trades off economic costs with operational risks for power dispatch under uncertainty. However, how to solve the RLD problem with provably near-optimal performance still remains an open problem. This paper presents a learning-based solution to this challenge. We first design a data-driven formulation for the RLD problem, which aims to construct a decision rule that directly maps day-ahead observable information to cost-effective dispatch decisions for the future delivery interval. Unlike most existing works that follow a predict-then-optimize paradigm, this end-to-end rule bypasses the additional suboptimality introduced by separately handling prediction and optimization. We then propose neural RLD, a novel solution method to the data-driven formulation. This method leverages an L2-regularized neural network to learn the decision rule, thereby transforming the data-driven formulation into a neural network training task that can be efficiently completed by stochastic gradient descent. A theoretical performance guarantee is further established to bound the suboptimality of our method, which implies that its suboptimality approaches to zero with high probability as more samples are utilized. Simulation tests across various systems demonstrate our method’s superior performance in convergence, suboptimality, and computational efficiency compared with benchmarks.
Index Terms:
Risk limiting dispatch, renewable integration, deep learning, stochastic programming, generalization performance guarantee.
I Introduction
The increasing adoption of uncertain renewables necessitates methodologies for power dispatch under uncertainty to achieve safe and cost-effective power system operations [1]. These methodologies typically fall into two major categories: i) robust optimization that safeguards the system performance for the worst case [2] and ii) stochastic programming and stochastic control formulations that optimize the expected performance [3].
Within the second category, Risk Limiting Dispatch (RLD) distinguishes itself as a class of methods that effectively trade off the economic cost with operation risks utilizing the most updated information about the uncertainties [4, 5]. Varaiya et al. [4] initially proposed RLD within a multi-stage stochastic control formulation to incorporate recourse decisions in multi-settlement markets. This approach was later generalized to explicitly leveraging forecast updates (i.e., the progressively improving forecasts for loads and renewable generation as we approach the delivery time) [5], generator ramping constraints [6], and energy storage [7]. These early studies mostly relied on deriving on analytical solutions for the underlying stochastic control problems, circumventing the curse of dimensionality associated with continuous state/action/disturbance variables in the original RLD formulation. However, when power network constraints are considered (e.g., as in the network RLD problem), multi-stage stochastic control formulation for RLD turns out to be intractable due to the high-dimensional continuous state space and the lack of analytical solutions. In fact, even the two-stage special case of the network RLD problem is challenging, for which Zhang et al. [8] proposed a heuristic algorithm and established its optimality under a “small-$\sigma$ assumption”, where $\sigma$ is the standard deviation of the forecast errors. It is noteworthy that the formulation and proposed algorithm in [8] do not rely on sampling from the probability distribution of the disturbances, which can be computationally costly when the disturbances are high-dimensional (e.g., considering correlated spatially disbursed wind and solar outputs). This makes their proposed algorithm easy to be incorporated in existing power system dispatch software (e.g., by modifying the reserve margin based on RLD solutions). Variants of RLD formulations have subsequently found a number of other applications in power system operations and planning [9, 10, 11]. Despite these prior studies, it remains an open problem to develop provably near-optimal methods that solve the RLD problem with network constraints and practical uncertainty levels (i.e., without the small-$\sigma$ assumption).
Another challenge associated with RLD and power dispatch under uncertainty in general is to obtain accurate characterizations for the uncertainties. Stochastic control formulations of RLD require the joint probability distribution of all uncertain parameters involved, which may be difficult to estimate from data when a large number of distributed wind and solar generators are considered [12]. Probabilistic forecasting methods based on machine learning [13, 14] could be leveraged to estimate these joint distributions. Other formulations have been proposed to address power dispatch under uncertainty, including scenario-based methods which represent the high-dimensional joint distributions using discrete scenarios [15, 16]. However, theoretical understanding of the performance of these formulations for characterizing uncertainty is lacking. Even when guarantees are available, it remains unclear how errors in quantifying the uncertainty may impact subsequent dispatch tasks derived from the uncertainty model.
Even if the difficulties associated with characterizing the uncertainties, i.e., the prediction (albeit probabilistic) step, and identifying the optimal decisions given the uncertainties, i.e., the optimization step, can be resolved, fundamental pitfalls around the two-step predict-then-optimize paradigm make it challenging to achieve direct data-driven, end-to-end optimal or provably near optimal RLD decisions. Indeed, it is common for the prediction step to use standard statistical loss functions, e.g., mean squared error (MSE), to quantify the accuracy ignoring the actual objective functions in the subsequent optimization step, and the optimization step to assumes the uncertainty model obtained from the prediction step is precise. However, when the errors in the prediction step is accounted for, a “more accurate” model for uncertainty as measured by MSE may not always lead to a better decision in the optimization step [17, 18]. Because of this “mismatch” in the objectives of the two steps, this paradigm may introduce additional suboptimality, resulting in undesirable performance.
I-A Contributions and paper organization
In this paper, we propose neural RLD, a formulation and method that merge the prediction and optimization steps, which directly use the data to train a neural network outputting dispatch decisions. This enables us to circumvent the challenges associated with the individual steps in the overall pitfalls of the predict-then-optimize paradigm, while obtaining end-to-end generalization bounds for the quality of the resulting dispatch decisions. This paper contributes to the literature in the following ways:
  1. 1.
    We design a data-driven formulation for the network RLD problem. This formulation aims to learn a decision rule from historical data. This decision rule is end-to-end, meaning it directly maps day-ahead observable information to cost-effective dispatch decisions for future delivery intervals, thereby bypassing the “mismatch” issue of the predict-then-optimize paradigm.
  2. 2.
    We propose the neural RLD method to solve the data-driven formulation above. This method utilizes an L2-regularized neural network to learn the decision rule, effectively converting the data-driven formulation into a neural network training task with a specialized training loss. We then explicitly derive the gradient of this specialized loss, enabling efficient completion of this training task by using stochastic gradient descent (SGD).
  3. 3.
    We develop a generalization performance guarantee based on statistical learning tools to theoretically bound the suboptimality of the neural RLD method. This guarantee implies that our method is asymptotically optimal, i.e., its suboptimality is likely to gradually decrease to zero with high probability as more samples are used.
The remaining parts are organized as follows. Section I-B reviews related literature. Section II describes the problem formulation of the network RLD. Section III introduces the proposed neural RLD. Section IV demonstrates simulation results, and Section V concludes this paper.
I-B Other related literature
I-B1 Learning-based stochastic dispatch
The application of machine learning to stochastic dispatch presents both innovative solutions and challenges. Initially, machine learning methods were usually deployed in one of two steps in the predict-then-optimize paradigm: enhancing prediction accuracy in the prediction step [19], or improving computational efficiency in the optimization step [20]. However, the aforementioned mismatch issue may bring additional suboptimality. In addressing this limitation, end-to-end frameworks, such as imitation learning [21], emerge as an alternative. By directly mapping observed features to dispatch decisions, imitation learning bypasses the mismatch issue. However, these end-to-end methods have been limited to deterministic problems [22]. Moreover, they often use the mean square error (MSE) between actual and predicted optimal decisions as the training loss. This choice might not accurately reflect the suboptimality of predicted decisions [17]. Additionally, despite demonstrating practical effectiveness in numerical case studies, these methods frequently fall short in providing theoretical performance assurances under varying conditions.
I-B2 Sample average approximation
Sample average approximation (SAA) is a method [23] that uses a finite number of discrete scenarios to approximate uncertainties, which can effectively transform RLD into a deterministic form. Due to its ease of use, it has been applied to various dispatch tasks, including economic dispatch [24], unit commitment [25], and optimal power flow [26]. Meanwhile, some error bounds have also been proposed to explain its performance on characterizing uncertainty. For example, Shapiro et al. [27] have shown that the SAA solution uniformly converges to the optimal one as the number of scenarios increases indefinitely. Additionally, Kim et al. [28] established a link between the number of scenarios used and the worst-case optimality gap of the SAA. However, the SAA still suffers from the aforementioned “mismatch” issue, as it is typically incorporated in the optimization step of the predict-then-optimize paradigm.
I-B3 Direct learning
In conference paper [12], we initially proposed an end-to-end learning method to predict RLD decisions directly from historical data. This method offered a generalization guarantee to bound its suboptimality. However, the initial study was limited to scalar decision variables, which significantly restricted its applicability and practicality in scenarios involving power network constraints. Meanwhile, it focused on theoretical analysis without numerical case studies. The current paper significantly extends the scope of [12] by concentrating on the network RLD with vectorial variables, providing a new generalization performance guarantee for this networked case, and conducting a series of case studies to validate the practical applicability of our method.
II Problem Formulation
We start by summarizing the classical network RLD formulation pertaining to the optimization step in the predict-then-optimize paradigm. We will then introduce a data-driven formulation of the problem that merges the prediction and optimization steps.
II-A Classic network risk limiting dispatch
As in [8], we consider a two-stage formulation of network RLD, where the system operator (SO) makes dispatch decisions for a future delivery interval sequentially in the day-ahead and real-time stages. Detailed discussion on how this formulation fits with the two-settlement electricity markets and the practical dispatch procedures can be found in [4, 8].
Consider a power network with ${N_{\mathrm{B}}}$ buses. In the day-ahead stage, the SO dispatches generators based on the most updated information regarding the net demand (i.e., load and renewable generation) over the network. Denote this dispatch decision by $\mathbf{u}\in\mathbb{R}^{N_{\mathrm{B}}}$, where $u_{n}$ is the generation at bus $n$. As the actual net demand $\mathbf{d}\in\mathbb{R}^{N_{\mathrm{B}}}$ is not available at day ahead, classical RLD formulation assumes that the probability distribution of $\mathbf{d}$ can be estimated. This probability distribution can depend on various auxiliary factors (e.g., temperature and wind speed information from the day-ahead weather forecast), and thus is denoted by $\mathcal{D}(\mathbf{x})$, where $\mathbf{x}\in\mathbb{R}^{p}$ is the $p\times 1$ vector of auxiliary information available at day ahead. In the real-time stage, the SO observes the realization of net demand $\mathbf{d}$ and can adjust the day-ahead dispatch by either dispatching extra generation or reducing the day-ahead dispatched amount, so the adjusted total generation profile balances the grid. We denote the real-time adjustment by $\mathbf{g}\in\mathbb{R}^{N_{\mathrm{B}}}$. Mathematically, the two-stage network RLD problem takes the form of
$\min_{\mathbf{u}}\quad\bm{\alpha}^{\intercal}\mathbf{u}+\mathbb{E}_{\mathbf{d}\sim\mathcal{D}(\mathbf{x})}[Q(\mathbf{u},\mathbf{d})],$ (1)
where $\bm{\alpha}\in\mathbb{R}^{N_{\mathrm{B}}}$ is the vector of linear generation cost coefficient for the day-ahead stage, and $Q(\mathbf{u},\mathbf{d})$ is the optimal second-stage cost given the day-ahead dispatch $\mathbf{u}$ and the realized net demand $\mathbf{d}$, defined as the optimal value of the real-time dispatch problem. In particular, as in [8], we use the following optimal DC (i.e., linearized AC) power flow formulation for the real-time dispatch:
$Q(\mathbf{u},\mathbf{d}):=\min_{\mathbf{g},\bm{\theta}}\quad$ $\bm{\beta}^{\intercal}(\mathbf{g})_{+},$ (2a)
s.t. $\mathbf{u}+\mathbf{g}-\mathbf{d}=\mathbf{B}\bm{\theta}\qquad\,:\bm{\lambda},$ (2b)
$-\mathbf{f}^{\mathrm{max}}\leq\mathbf{F}\bm{\theta}\leq\mathbf{f}^{\mathrm{max}}:\bm{\nu}_{-},\bm{\nu}_{+},$ (2c)
where the objective function models the cost of dispatching extra real-time generation or shedding loads with $(\mathbf{g})_{+}:=\max\{\mathbf{g},\mathbf{0}\}$ and $\bm{\beta}\in\mathbb{R}^{N_{\mathrm{B}}}$ modeling the per unit cost of extra generation or value of loss load. Eq. (2b) represents the DC power flow equation, where matrix $\mathbf{B}\in\mathbb{R}^{{N_{\mathrm{B}}}\times{N_{\mathrm{B}}}}$ maps voltage phase angles $\bm{\theta}\in\mathbb{R}^{N_{\mathrm{B}}}$ into nodal active power injections. Eq. (2c) denotes the branch capacity constraints, where $\mathbf{F}\in\mathbb{R}^{{N_{\mathrm{L}}}\times{N_{\mathrm{B}}}}$ maps $\bm{\theta}$ to branch flows, $\mathbf{f}^{\mathrm{max}}$ is the corresponding maximum allowable values, and ${N_{\mathrm{L}}}$ is the number of branches. Vectors $\bm{\lambda}\in\mathbb{R}^{N_{\mathrm{B}}}$, $\bm{\nu}_{-}\in\mathbb{R}^{N_{\mathrm{L}}}$ and $\bm{\nu}_{+}\in\mathbb{R}^{N_{\mathrm{L}}}$ are the dual variables of (2b)-(2c), respectively.
The network RLD problem is typically tackled by a two-step predict-then-optimize paradigm, i.e., first estimating the distribution $\mathcal{D}$ via probabilistic forecasting, and then solving the stochastic problem (1). However, as mentioned in Section I, separately considering the prediction and optimization leads to challenges for each step and loss of optimality overall.
II-B Data-driven risk limiting dispatch
We next introduce a new formulation of two-stage network RLD that merges the prediction and optimization steps in the predict-then-optimize paradigm. In other words, we will leverage the historical data that are commonly used in the prediction step to directly learn a decision rule for the future delivery interval. We refer to this formulation as data-driven network RLD.
The historical dataset includes day-ahead auxiliary information $\mathbf{x}^{(m)}$ and the realized net demand $\mathbf{d}^{(m)}$ for a collection of past delivery intervals $m\in\mathcal{M}$ with $M:=|\mathcal{M}|$. As this dataset $\mathcal{S}:=\{(\mathbf{x}^{(m)},\mathbf{d}^{(m)}),m\in\mathcal{M}\}$ will be used to train a decision rule for future delivery intervals, we refer to $\mathcal{S}$ as the training set. The training set is usually available to the SO well in advance of the delivery interval for which the SO needs to make dispatch decisions about. At the day-ahead stage for the delivery interval of concern, the SO will also have access to the auxiliary information for the particular delivery interval, denoted by $\mathbf{x}^{\prime}$. In the prediction step of the conventional predict-then-optimize paradigm, if a deterministic forecast is used in place of a probabilistic one for the future net demand, $\mathbf{x}^{\prime}$ will be used together with a machine learning model trained on $\mathcal{S}$ for demand forecast. In our setting, we wish to apply the decision rule trained from $\mathcal{S}$ at $\mathbf{x}^{\prime}$ to predict the RLD decisions. We refer to $\mathbf{x}^{\prime}$ as the test input.
Ideally, we wish to find the optimal decision rule $\mathbf{h}^{\star}:\mathbb{R}^{p}\mapsto\mathbb{R}^{N_{\mathrm{B}}}$, that maps the day-ahead auxiliary information $\mathbf{x}$ to the optimal dispatch decision $\mathbf{u}^{\star}$. [^1] Given $\mathbf{u}$ and the realization of $\mathbf{d}$, it is straightforward to calculate the optimal real-time dispatch by solving (II-A). Thus it suffices to focus on the day-ahead decision $\mathbf{u}$. Adapting terminologies from the statistical learning theory, we will also refer to such a decision rule as a hypothesis. Mathematically, supposing that there is an underlying true joint distribution of $(\mathbf{x},\mathbf{d})$, denoted by $\mathcal{Z}$, we can write the following problem to identify the optimal hypothesis as
$\mathbf{h}^{\star}=\mathop{\arg\min}_{\mathbf{h}\in\mathcal{H}}\ \mathbb{E}_{(\mathbf{x},\mathbf{d})\sim\mathcal{Z}}\left[\bm{\alpha}^{\intercal}\mathbf{h}(\mathbf{x})+Q(\mathbf{h}(\mathbf{x}),\mathbf{d})\right],$ (3)
where $\mathcal{H}$ is a hypothesis class containing all functions that map $\mathbf{x}$ to $\mathbf{u}$. However, directly evaluating the expectation in (3) is challenging as the distribution $\mathcal{Z}$ is usually not available. Instead, we approximate the expectation with its empirical mean using the training set $\mathcal{S}$ as done in empirical risk minimization (ERM) [29]:
$\widehat{\mathbf{h}}=\mathop{\arg\min}_{\mathbf{h}\in\mathcal{H}}\frac{1}{M}\sum_{m\in\mathcal{M}}\left(\bm{\alpha}^{\intercal}\mathbf{h}(\mathbf{x}^{(m)})+Q(\mathbf{h}(\mathbf{x}^{(m)}),\mathbf{d}^{(m)})\right).$ (4)
We refer to (4) as the data-driven RLD problem. Upon solving (4), and at the day-ahead stage for the future delivery time of concern, we can directly obtain the dispatch decision via $\mathbf{u}^{\prime}=\widehat{\mathbf{h}}(\mathbf{x}^{\prime})$.
This data-driven formulation bypasses the “mismatch” issue discussed previously by merging the prediction and optimization steps. Moreover, it does not need data of the optimal decision $\mathbf{u}^{\star}$ for historical delivery interval $m\in\mathcal{M}$ as the training labels. However, it poses challenges for two primary reasons. First, given that $\mathbf{h}\in\mathcal{H}$ can represent any $\mathbb{R}^{N_{\mathrm{B}}}$-valued function, Eq. (4) is computationally intractable due to its infinite-dimensional nature. Second, we only have a finite number of historical data points in the training set. These data points may fail to accurately reflect the true distribution of $\mathcal{Z}$, so the ERM hypothesis $\widehat{\mathbf{h}}$ may perform much worse than $\mathbf{h}^{\star}$.
III Neural Risk Limiting Dispatch
Motivated by the above challenges, we propose the neural RLD. In this approach, the hypothesis class $\mathcal{H}$ is represented by L2 regularized neural networks with a finite number of parameters. Consequently, problem (4) is redefined as a finite-dimensional neural network training task with a specialized training loss. To support this approach, we provide a generalization performance guarantee based on statistical learning tools including uniform convergence and Rademacher complexity, which theoretically bounds the optimality gap of our method with high probability. [^2] It is impossible to obtain performance guarantees here with probability one as there is always a non-zero probability (as long as the training set is finite) to obtain a training set for which the empirical mean in (4) is not a good representation of the population mean in (3). This guarantee demonstrates that as more historical data points are utilized, the performance of our neural RLD closely approximates that of the optimal hypothesis in the same hypothesis class. In this section, we first describe the overall neural RLD procedure. Following this, we delve into the training process of the neural RLD. Finally, we establish the generalization performance guarantee.
III-A Neural network for data-driven RLD
The proposed neural RLD chooses L2-regularized neural networks as the hypothesis class due to their excellent approximation capability [30]. By using $k\in\mathcal{K}$ and $j\in\mathcal{J}^{k}$ with $K:=|\mathcal{K}|$ and $J^{k}:=|\mathcal{J}^{k}|$ to index the hidden layers and the neurons in each, this hypothesis class can be expressed as:
=$\mathcal{H}=\left\{\mathbf{h}(\mathbf{x})\left|\begin{aligned} &\mathbf{h}(\mathbf{x})=\mathbf{W}^{K}\phi\left(\cdots\mathbf{W}^{2}\phi\left(\mathbf{W}^{1}\mathbf{x}\right)\right),\\ &\bm{\|}\mathbf{w}^{k}_{j}\|_{2}\leq W^{\max},\forall j\in\mathcal{J}^{k},\forall k\in\mathcal{K},\end{aligned}\right.\right\},$ (5)
where $\mathbf{h}:\mathbb{R}^{p}\mapsto\mathbb{R}^{N_{\mathrm{B}}}$ represents a neural network in this hypothesis class; $\mathbf{W}^{k}\in\mathbb{R}^{J^{k+1}\times J^{k}}$ is the weights of the $k$-th layer and $\mathbf{w}_{j}^{k}\in\mathbb{R}^{J^{k}}$ is the transpose of its $j$-th row. Symbol $\phi(\cdot)$ denotes the activation function. This paper chooses ReLU as the activation function. It operates on each individual element of the input, ensuring that the dimensions of both input and output remain identical. The second constraint in (5) regularize the class of neural networks by limiting the L2 norm of the weights, where $W^{\max}$ is the maximum norm of weights allowed. If we further define a new loss function $\ell$, also referred as to the RLD loss, as follows:
$\ell(\mathbf{h}(\mathbf{x}),\mathbf{d})=\bm{\alpha}^{\intercal}\mathbf{h}(\mathbf{x})+Q(\mathbf{h}(\mathbf{x}),\mathbf{d}),$ (6)
then problem (4) can be interpreted as a neural network training task. This task uses $\mathcal{S}=\{(\mathbf{x}^{(m)},\mathbf{d}^{(m)}),\forall m\in\mathcal{M}\}$ as the training set, and leverages $\ell$ as the training loss, as follows:
$\widehat{\mathbf{h}}=\mathop{\arg\min}\limits_{h\in\mathcal{H}}\frac{1}{M}\sum_{m\in\mathcal{M}}\ell(\mathbf{h}(\mathbf{x}^{(m)}),\mathbf{d}^{(m)}).$ (7)
III-B Neural network training
The training task (7) can be effectively completed using the stochastic gradient descent (SGD) algorithm if the (sub-) gradient of the RLD loss (6) with respect to the weights exists and can be effectively evaluated. Based on the chain rule, this gradient is expressed as:
$\nabla_{\mathbf{w}_{j}^{k}}\ell=\left(\bm{\alpha}+\nabla_{\mathbf{u}}Q\right)^{\intercal}\nabla_{\mathbf{w}_{j}^{k}}\mathbf{u},\forall j\in\mathcal{J}^{k+1},\forall k\in\mathcal{K},$ (8)
where $\mathbf{u}\in\mathbb{R}^{{N_{\mathrm{B}}}}$ is the output of the hypothesis $\mathbf{h}$, and we have $\nabla_{\mathbf{w}_{j}^{k}}\ell\in\mathbb{R}^{J^{k}}$, $\nabla_{\mathbf{u}}Q\in\mathbb{R}^{{N_{\mathrm{B}}}}$, and $\nabla_{\mathbf{w}_{j}^{k}}\mathbf{u}\in\mathbb{R}^{{N_{\mathrm{B}}}\times J^{k}}$. The gradient $\nabla_{\mathbf{w}_{j}^{k}}\mathbf{u}$ can be easily obtained using automatic differentiation in modern deep learning frameworks like Pytorch. However, the gradient $\nabla_{\mathbf{u}}Q$ is hard to calculate because $Q(\mathbf{u},\mathbf{d})$ involves an optimization problem (II-A). To address this issue, we leverage the strong duality to explicitly express $\nabla_{\mathbf{u}}Q$, as described in the following proposition.
Proposition 1 (Expression of the gradient).
The gradient $\nabla_{\mathbf{u}}Q$ can be expressed as:
$\nabla_{\mathbf{u}}Q=\bm{\lambda}^{\star},$ (9)
where $\bm{\lambda}^{\star}$ is the optimal dual variable of constraint (2b).
Proof: See Appendix A.
Then, we can introduce the SGD to train the neural network. The whole training process is summarized in Algorithm 1. Here, we use $N^{\mathrm{ep}}$ and $\mathcal{B}$ to denote the maximum epoch number and the sample index of a batch, respectively.
Input : Dataset $\mathcal{S}=\{(\mathbf{x}^{(m)},\mathbf{d}^{(m)}),\forall m\in\mathcal{M}\}$, learning rate $\rho$, and batch size $B=|\mathcal{B}|$
Output : The hypothesis ${\mathbf{h}}$
1 Initialize weights $\mathbf{W}^{k},\forall k\in\mathcal{K}$;
2 for $epoch\leftarrow 1$ to $N^{\mathrm{ep}}$ do
3       foreach batch $\{(\mathbf{x}^{(m)},\mathbf{d}^{(m)}),\forall{m}\in\mathcal{B}\}$ do
4            ${\mathbf{u}}^{(m)}\leftarrow\mathbf{h}({\mathbf{x}}^{(m)}),\forall m\in\mathcal{B}$;
5             Solve the dual of (II-A) to get $(\bm{\lambda}^{\star})^{(m)},\forall{m}\in\mathcal{B}$;
6             $\mathbf{w}_{j}^{k}\leftarrow\mathbf{w}_{j}^{k}-\frac{\rho}{B}\sum_{m\in\mathcal{B}}\left(\bm{\alpha}+(\bm{\lambda}^{\star})^{(m)}\right)^{\intercal}\nabla_{\mathbf{w}_{j}^{k}}\mathbf{u}^{(m)},\ \forall j\in\mathcal{J}^{k},\forall k\in\mathcal{K}$;
7
8       end foreach
9
10 end for
Output ${\mathbf{h}}$.
Algorithm 1 Neural network training
III-C Generalization performance guarantee
This section provides a theoretical generalization performance guarantee, which bounds the optimality gap of the proposed neural RLD with a large probability. Specifically, we first define the generalization cost of a hypothesis $\mathbf{h}$, denoted by $L(\mathbf{h})$, as follows:
$L(\mathbf{h})=\mathbb{E}_{(\mathbf{x},\mathbf{d})\sim\mathcal{Z}}\left[\ell(\mathbf{h}(\mathbf{x}),\mathbf{d})\right].$ (10)
This generalization cost is the same as the objective in (3), which represents the expected cost of the decision given by $\mathbf{h}$ over the true (and unknown) population distribution $\mathcal{Z}$. For any hypothesis $\mathbf{h}$, we are also interested how its generalization cost compared to that of the optimal one $\mathbf{h}^{\star}$:
$\Delta L(\mathbf{h})={L}(\mathbf{h})-{L}(\mathbf{h}^{\star}),$ (11)
where $\Delta L(\mathbf{h})$ is referred to as the excess cost of $\mathbf{h}$.
As introduced in Section III-A, we train a L2-regularized neural network $\widehat{\mathbf{h}}$ to act as our hypothesis. In order to calculate the excess cost of $\widehat{\mathbf{h}}$, two mild assumptions are in force:
  • A1: Data points in dataset $\mathcal{S}$ are identically and independently distributed (i.i.d.), and follow an unknown distribution $\mathcal{Z}$.
  • A2: The expectation of day-ahead auxiliary information vector $\mathbf{x}$ is bounded, i.e., $\mathbb{E}_{\mathbf{x}}\left(\|\mathbf{x}\|_{2}^{2}\right)\leq(X^{\mathrm{max}})^{2}$ for some finite $X^{\max}$.
While Assumption A1 is commonly used in the prediction step of the predict-then-optimize paradigm (e.g., in probabilistic forecasting settings), especially when theoretical performance guarantees are developed, the precise form of it may not be satisfied in many application settings. This does not fundamentally limit the applicability of our results for two reasons: (a) In any applications where past data is used to guide future decisions, some forms of stationarity assumption is necessary (so past data indeed contain relevant information for the decision). Our results can be extended to settings with weaker forms of stationarity assumption leveraging statistical learning theory tools for non-i.i.d. data [31]. (b) When the historical dataset is not stationary as it is, there may be transformations (e.g., de-trending and differencing as done in time series analysis) to remove non-stationarity.
We next derive a probably approximately correct (PAC) upper bound for the excess cost of $\widehat{\mathbf{h}}$, i.e., a bound for the suboptimality of $\widehat{\mathbf{h}}$ that holds with high probability. This is achieved by applying statistical learning theory tools (i.e., uniform convergence) to our neural RLD hypothesis class and loss function, for which we develop our own Lipschitz continuity bound and Rademacher complexity bound, [^3] Rademacher complexity measures the richness of a hypothesis class. A hypothesis class with a high Rademacher complexity in general is more expressive but requires more data points to learn (e.g., with more parameters). respectively. The final result is stated as follows:
Theorem 1 (PAC bound of excess cost).
When Assumptions A1 and A2 hold, with probability at least $1-\delta$ for any small $\delta\in(0,1)$, the excess cost of hypothesis $\widehat{\mathbf{h}}$ is bounded by:
$\Delta L(\widehat{\mathbf{h}})\leq\frac{4(2{W^{\max}})^{K-\frac{1}{2}}C_{\ell}X^{\mathrm{max}}+\sqrt{2\ln(2/\delta)}}{\sqrt{M}}.$ (12)
where $C_{\ell}$ is a finite constant related to the Lipschitz continuity of the RLD loss $\ell$.
Proof: Appendix B provides the detailed steps for deriving (12) and the expression for $C_{\ell}$.
By (11), the PAC bound provided by Theorem 1 can be regarded as the expected optimality gap of the hypothesis $\widehat{\mathbf{h}}$. Thus, we also refer to Theorem 1 as the generalization performance guarantee of hypothesis $\widehat{\mathbf{h}}$. This PAC bound decreases to zero as more samples are used, which indicates that the performance of $\widehat{\mathbf{h}}$ gradually approaches to the optimal hypothesis $\mathbf{h}^{\star}$ with probability $1-\delta$ even for arbitrarily small $\delta>0$. Hence, we have the following colloary.
Colloary 1 (Asymptotic optimality).
The neural RLD hypothesis $\widehat{\mathbf{h}}$ is asymptotically optimal, i.e., with probability at least $1-\delta$, $\Delta L(\widehat{\mathbf{h}})\to 0$ as $M\to\infty$ for any $\delta\in(0,1)$.
IV Case Study
IV-A Simulation setup
IV-A1 Test systems
To validate the performance of the proposed framework, we implement our case studies based on four different systems, i.e, the IEEE 5-bus, 118-bus, 300-bus, and 1354-bus test systems. The numbers of buses, line, loads, and generators in different test systems are summarized in Table I. The unit costs in the first-stage and second-stage, i.e., $\bm{\alpha}$ and $\bm{\beta}$, are illustrated in Fig. 1. Other parameters remain the same as the standard IEEE test cases.
TABLE I: Descriptions of our test cases and structures of the corresponding neural networks
Cases Line number Load number Generator number Structures of of neural networks
5-bus 6 5 5 (5, 5, 5)
118-bus 173 99 54 (20, 20, 20)
300-bus 411 199 69 (40, 40, 40)
1354-bus 1991 673 260 (60, 60, 60)
(a) 5-bus case
(b) 118-bus case
(c) 300-bus case
(d) 1354-bus case
Figure 1: Unit generation costs used in different test cases. Note the first-stage unit generation cost $\bm{\alpha}$ is always smaller than the second-stage one $\bm{\beta}$.
IV-A2 Benchmarks
The following two benchmarks are introduced to verify the benefits of the proposed neural RLD:
  • Two-step predict-then-optimize paradigm used in [32] (termed “Two-step”). It first follows [14] and trains a mixture density network to learn the distribution of net demands. According to our test, this model shows high accuracy, e.g., the relative errors for predicting the mean and covariance are smaller than 3%. Then, the scenario-based method is employed to solve problem (II-A), in which 30 discrete scenarios are generated to approximate the predicted distribution.
  • Imitation learning used in [21] (termed “Imitation”). This benchmark trains a neural network using the MSE loss to directly learn the mapping from auxiliary information $\mathbf{x}$ to the dispatch decision $\mathbf{u}$. The training labels, i.e., historical optimal decisions, are generated by solving the deterministic dispatch problems given the realizations of $\mathbf{d}$ in the past delivery intervals.
IV-A3 Data generation
For each test case, we randomly generate 5,000 samples of $\mathbf{x}\in\mathbb{R}^{p}$ from a uniform distribution on the interval $[0,1]^{p}$. Then, the uncertain net demand $\mathbf{d}\in\mathbb{R}^{{N_{\mathrm{B}}}}$ is generated based on the following rule: [^4] One may anticipate good prediction performance and perhaps superior performance of the Two-step paradigm due to the linearity between $\mathbf{d}$ and $\mathbf{x}$ in data generation. While the former observation is indeed the case, the latter turned out to be false due to the mismatch of the two steps in the predict-then-optimize paradigm. In fact, we choose the simple rule (13) precisely to highlight the importance of end-to-end decision making.
$\mathbf{d}=\bm{\Omega}\left(\mathbf{x}*(\bm{1}+0.15\cdot\bm{\pi})\right),$ (13)
where $\bm{\Omega}\in\mathbb{R}^{{N_{\mathrm{B}}}\times p}$ is a fixed coefficient matrix; *** represents the element-wise multiplication. A random variable $\bm{\pi}\in\mathbb{R}^{p}$ is introduced to simulate the stochastic characteristic of net demands. Here, we assume that $\bm{\pi}$ follows the standard Gaussian distribution. 4,000 samples are treated as the training set, while the rest ones are used as the testing set to verify the suboptimality of different methods.
IV-A4 Structures of neural networks
The neural networks used in all methods are composed of fully connected layers. Their structures are summarized in the column “Structures of neural networks” in Table I, where every number in brackets represents the neuron number in a hidden layer. For example, (5, 5, 5) represents the corresponding neural network has three hidden layers, each with five neurons. Note the neural networks maintain the same structures among all methods.
IV-B Performance evaluation
IV-B1 Convergence
Fig. 2 demonstrates the training losses of the neural networks in each test case. Note all the training losses are standardized to [0,1] for a fair comparison. In every case, the proposed method can converge in only 100 epochs. In contrast, both the benchmarks may not converge even after 1,000 epochs. According to Proposition 1, the gradient of the RLD loss is well-defined, so the proposed model can be efficiently trained by the SGD. In benchmark Imitation, the neural network is trained using the MSE to learn the mapping from auxiliary information $\mathbf{x}$ to the optimal decision $\mathbf{u}^{\star}$. Since the RLD problem (1) is not strictly convex, this mapping may be discontinuous [17], which is hard to learn. Benchmark Two-step trains a mixture density network to learn the distribution of net demands for a given $\mathbf{x}$ [14]. This mixture density network needs to predict the means and covariances for multiple Gaussian components, while a set of parameters need to be learned for each Gaussian component. Thus, it has more learnable parameters than the common neuron network even if they have the same hidden layers and neurons, leading to lower training efficiency.
(a) 5-bus case
(b) 118-bus case
(c) 300-bus case
(d) 1354-bus case
Figure 2: Standardized training losses of the neural networks trained by different methods. Note the horizontal axis is logarithmic.
IV-B2 Suboptimality
We compare the normalized suboptimality of different methods, which is defined as:
$\mathrm{Suboptimality}=\frac{\ell(\widehat{\mathbf{u}},\mathbf{d})-\ell({\mathbf{u}^{\star}},\mathbf{d})}{\ell({\mathbf{u}^{\star}},\mathbf{d})},$ (14)
where $\widehat{\mathbf{u}}$ represents the dispatch decision given by each method. The numerator represents the excess generation cost between $\widehat{\mathbf{u}}$ and the optimal decision ${\mathbf{u}^{\star}}$, while the denominator is the optimal objective of (II-A). We calculate this normalized suboptimality for every data point in the testing set. The results are given in Fig. 3. In all test cases, the normalized suboptimality of the proposed method is significant smaller than those of the two benchmarks. For instance, the average suboptimality of the proposed method is only 1.46% in the 300-bus test case, while this value is 5.54% and 3.07% in benchmarks Imitation and Two-step, respectively. As aforementioned, the RLD loss can more accurately reflect the suboptimality of decisions compared to the MSE loss used in benchmark Imitation. Moreover, the proposed neural network is an end-to-end optimization method and can bypass the “mismatch” issue of benchmarks Two-step. Thus, it exhibits the lowest suboptimality among all methods.
(a) 5-bus case
(b) 118-bus case
(c) 300-bus case
(d) 1354-bus case
Figure 3: Normalized suboptimality given by different methods.
IV-B3 Computational efficiency
Fig. 4 compares the required time for solving one single instance and 5000 instances of (1) using different methods, where benchmark Two-step is solved in parallel with 30 threads. Since both the proposed neural RLD and benchmark Imitation replace the solving process with the forward pass of neural networks, they can output dispatch decisions in a very short time. On the contrary, benchmark Two-step still needs to solve (1), so it requires much longer time. For example, the solving times required by the proposed method and benchmark Imitation is at least two orders of magnitude smaller than that of benchmark Two-step.
(a) 5-bus case
(b) 118-bus case
(c) 300-bus case
(d) 1354-bus case
Figure 4: Times for solving one and 5,000 instances with different methods.
IV-C Sensitivity analysis
IV-C1 Neuron number
We investigate the impacts of neuron numbers on the normalized suboptimality of different methods based on the 118-bus test system. The results are illustrated in Fig. 5. The proposed method consistently exhibits the lowest suboptimality in all cases. Meanwhile, even with only a few neurons, the proposed method demonstrates low suboptimality. This superiority results from the RLD loss, enabling more accurate evaluation on the suboptimality of decisions compared to the MSE. This conclusion can be further validated by Fig. 6, in which the MSE values and average suboptimality of benchmark Imitation method are illustrated. While increasing neuron number can reduce the MSE, the suboptimality does not exhibit a consistent decrease.
Figure 5: Normalized suboptimality of all methods under different neuron numbers in the 118-bus test case.
Figure 6: Training losses, i.e., the MSE, and average normalized suboptimality of the imitation learning under different neuron numbers. Obviously, a smaller MSE may not always result in lower suboptimality.
IV-C2 Sample size
Fig. 7 compares the normalized suboptimality of all methods with different sample sizes. Similar to the results in Fig. 6, the proposed method always outperforms the rest ones by achieving the lowest normalized suboptimality. Moreover, with an increasing sample size, the suboptimality of the proposed method decreases. According to our proposed generalization performance guarantee, i.e., Theorem 1, the PAC bound of our method gradually approaches zero with the increase of the sample size. Other methods do not display a similar trend in suboptimality.
Figure 7: Normalized suboptimality of all methods with different training sample sizes in the 118-bus test case.
V Conclusions
This paper presents a learning-based approach to solve the network RLD problem with provably near-optimal performance. It begins by designing a data-driven formulation for the network RLD problem, which aims to learn a decision rule from historical data. This decision rule directly maps day-ahead auxiliary information to cost-effective dispatch decisions for the future delivery interval, so it can bypass the inherent “mismatch” issue of the widely adopted predict-then-optimize paradigm. We then propose neural RLD as a novel solution method for this data-driven formulation. It employs a L2-regularized neural network to learn the decision rule, thereby transforming the above data-driven formulation as a neural network training task with a specialized training loss. The gradient of this loss is explicitly derived, allowing for efficient completion of the training task using SGD. A generalization performance guarantee is further developed based on uniform convergence, which provides a PAC upper bound for the suboptimality of the neural RLD. Our numerical experiments on IEEE 5-bus, 118-bus, 300-bus, and 1354-bus test systems demonstrate the neural RLD’s desirable performance in terms of convergence, suboptimality, and computational efficiency. Simulations also confirm the asymptotic optimality of our method, implying that its suboptimality diminishes to zero with high probability as more samples are utilized.
The proposed neural RLD offers an effective solution for the two-stage network RLD problem. However, its applicability to the original multi-stage stochastic control formulation remains limited. Additionally, our theoretical guarantees are currently confined to neural networks composed solely of fully connected layers. In view of these limitations, our future work aims to extend the neural RLD to the multi-stage formulation and to establish theoretical guarantees for more sophisticated learning models.
Appendix A
Proof of Proposition 1: The optimal objective of the dual problem of (II-A) is expressed as:
$V(\mathbf{u},\mathbf{d})=$ $\left(\bm{\lambda}^{\star}\right)^{\intercal}\mathbf{u}-(\bm{\lambda}^{\star})^{\intercal}\mathbf{d}-\left(\bm{\nu}_{-}^{\star}+\bm{\nu}_{+}^{\star}\right)^{\intercal}\mathbf{f}^{\mathrm{max}}.$ (15)
Since (II-A) can be reformulated as a linear program, the strong duality holds, so
$V(\mathbf{u},\mathbf{d})=Q(\mathbf{u},\mathbf{d}).$ (16)
Then, the derivative $\nabla_{\mathbf{u}}Q$ can be expressed as (9).
Appendix B
Proof of Theorem 1: This proof is composed of the following steps: 1) prove the Lipschitz continuity of the RLD loss (6); 2) obtain an upper bound of the Rademacher complexity, and 3) derive the PAC upper bound given in Theorem 1.
B-1 Lipschitz continuity of the RLD loss
Suppose $(\mathbf{g}^{\star})^{(1)}$ and $(\mathbf{g}^{\star})^{(2)}$ are the optimal solutions of (II-A) when $\mathbf{u}=\mathbf{u}^{(1)}$ and $\mathbf{u}=\mathbf{u}^{(2)}$, respectively. According to [Theorem 2.2] in [33], the following inequality holds because (II-A) can be transformed into a linear problem by introducing auxiliary variables:
$\|(\mathbf{g}^{\star})^{(1)}-(\mathbf{g}^{\star})^{(2)}\|_{\infty}\leq C_{\mathbf{g}}\|\mathbf{u}^{(1)}-\mathbf{u}^{(2)}\|_{2},$ (17)
where $C_{\mathbf{g}}$ is a finite constant. Then, we have:
$\left|Q(\mathbf{u}^{(1)},\mathbf{d}^{(1)})-Q(\mathbf{u}^{(2)},\mathbf{d}^{(2)})\right|$
$=$ $\left|\bm{\beta}^{\intercal}\left((\mathbf{g}^{\star})^{(1)}\right)_{+}-\bm{\beta}^{\intercal}\left((\mathbf{g}^{\star})^{(2)}\right)_{+}\right|$
$\leq$ $\|\bm{\beta}\|_{2}\cdot\|(\mathbf{g}^{\star})^{(1)}-(\mathbf{g}^{\star})^{(2)}\|_{2}$
$\leq$ $\sqrt{{N_{\mathrm{B}}}}\cdot\|\bm{\beta}\|_{2}\cdot\|(\mathbf{g}^{\star})^{(1)}-(\mathbf{g}^{\star})^{(2)}\|_{\infty}$
$\leq$ $\sqrt{{N_{\mathrm{B}}}}C_{\mathbf{g}}\cdot\|\bm{\beta}\|_{2}\cdot\|\mathbf{u}^{(1)}-\mathbf{u}^{(2)}\|_{2},$ (18)
where the first inequality holds because function $\mathbf{g}\rightarrow\bm{\beta}^{\intercal}(\mathbf{g})_{+}$ is $\|\bm{\beta}\|_{2}$-Lipschitz continuous; the second inequality holds by the relation between infinite and L2 norms. Finally, the Lipschitz continuity of $\ell(\mathbf{u},\mathbf{d})$ can be proved by:
$\left|\ell(\mathbf{u}^{(1)},\mathbf{d}^{(1)})-\ell(\mathbf{u}^{(2)},\mathbf{d}^{(2)})\right|$
$=$ $\left|\bm{\alpha}^{\intercal}\mathbf{u}^{(1)}+Q(\mathbf{u}^{(1)},\mathbf{d}^{(1)})-\bm{\alpha}^{\intercal}\mathbf{u}^{(2)}-Q(\mathbf{u}^{(2)},\mathbf{d}^{(2)})\right|$
$\leq$ $\underbrace{\left(\|\bm{\alpha}\|_{2}+\sqrt{{N_{\mathrm{B}}}}C_{\mathbf{g}}\|\bm{\beta}\|_{2}\right)}_{C_{\ell}}\|\mathbf{u}^{(1)}-\mathbf{u}^{(2)}\|_{2}.$ (19)
Note (19) also gives the definition of $C_{\ell}$ in Theorem 1. When the L2-regularized neural network $\mathbf{h}$ is used to predict the dispatch decision $\mathbf{u}$, we have:
$\|\mathbf{u}^{(1)}-\mathbf{u}^{(2)}\|_{2}$ $=\|\mathbf{h}(\mathbf{x}^{(1)})-\mathbf{h}(\mathbf{x}^{(2)})\|_{2}.$ (20)
By substituting (20) into (19), we obtain:
$\left|\ell(\mathbf{u}^{(1)},\mathbf{d}^{(1)})-\ell(\mathbf{u}^{(2)},\mathbf{d}^{(2)})\right|\leq C_{\ell}\|\mathbf{h}(\mathbf{x}^{(1)})-\mathbf{h}(\mathbf{x}^{(2)})\|_{2}.$ (21)
The above inequality shows that the mapping $\psi(\cdot):\mathbf{h}(\mathbf{x})\mapsto\ell(\mathbf{h}(\mathbf{x}),\mathbf{d})$ is $C_{\ell}$-Lipschitz continuous.
B-2 Upper bound of Rademacher complexity
We first define the following new hypothesis class for convenience:
$\mathcal{H}^{k}=\left\{\mathbf{h}^{k}(\mathbf{x})\left|\begin{aligned} &\mathbf{h}^{k}(\mathbf{x})=\mathbf{W}^{k}\phi\left(\cdots\mathbf{W}^{2}\phi\left(\mathbf{W}^{1}\mathbf{x}\right)\right),\\ &\bm{\|}\mathbf{w}^{k^{\prime}}_{j}\|_{2}\leq W^{\max},\forall j\in\mathcal{J}^{k^{\prime}},\\ &\quad\quad\quad\quad\quad\quad\ \ \ \ \forall k^{\prime}\in[1,\cdots,k],\end{aligned}\right.\right\},$ (22)
where the hypothesis $\mathbf{h}^{k}(\mathbf{x}):\mathbb{R}^{p}\mapsto\mathbb{R}^{J^{k}}$ represents a function maps $\mathbf{x}$ to the output of the $k$-th hidden layer.
The empirical Rademancher complexity of the loss class on dataset $\mathcal{S}$ is defined as follows:
^$\widehat{\mathcal{R}}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})=\mathbb{E}_{\xi}\left[\sup_{\mathbf{\ell}\in\mathcal{L}_{\mathbf{h}}}\sum_{m\in\mathcal{M}}\xi^{(m)}\ell(\mathbf{h}(\mathbf{x}^{(m)}),\mathbf{d}^{(m)})\right].$ (23)
where ^$\widehat{\mathcal{R}}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})$ is calculated based on observed realizations, which differs from the Rademancher complexity $\mathcal{R}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})$. To obtain the upper bound of ^$\widehat{\mathcal{R}}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})$, we first introduce the vector contraction inequality in the following lemma:
Lemma 1 (Vector contraction inequality [33]).
Let $\mathcal{H}$ be a class of vector-valued functions $\mathbf{h}(\mathbf{x}):\mathbb{R}^{p}\mapsto\mathbb{R}^{{N_{\mathrm{B}}}}$ and let $\psi(\cdot):\mathbb{R}^{{N_{\mathrm{B}}}}\mapsto\mathbb{R}$ is a $C_{\ell}$-Lipschitz continuous function. Then, the following vector contraction inequality holds:
$\mathbb{E}_{\xi}\left(\sup_{\mathbf{h}\in\mathcal{H}}\sum_{m\in\mathcal{M}}\xi^{(m)}\psi(\mathbf{h}(\mathbf{x}^{(m)}))\right)$
$\quad\ \leq\sqrt{2}C_{\ell}\mathbb{E}_{\xi}\left(\sup_{h_{j}\in\mathcal{H}_{j}}\sum_{m\in\mathcal{M}}\sum_{j=1}^{{N_{\mathrm{B}}}}\xi^{(m,j)}h_{j}(\mathbf{x}^{(m)})\right),$ (24)
where $\xi^{(m,j)}$ is a random variable and drawn i.i.d.. from a uniform distribution over $\{-1,+1\}$; $h_{j}(\mathbf{x}):\mathbb{R}^{p}\mapsto\mathbb{R}$ is a mapping from $\mathbf{x}$ to the $j$-th component of $\mathbf{h}$; $\mathcal{H}_{j}$ represents the hypothesis class of $h_{j}(\mathbf{x})$ defined such that $\mathbf{h}_{j}\in\mathcal{H}_{j}$.
By substituting Lemma 1 into (23), we have:
^$\widehat{\mathcal{R}}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})\leq\sqrt{2}C_{\ell}\mathbb{E}_{\xi}\left[\sup_{{h}_{j}\in\mathcal{H}_{j}}\sum_{m\in\mathcal{M}}\sum_{j=1}^{{N_{\mathrm{B}}}}\xi^{(m,j)}{h}_{j}(\mathbf{x}^{(m)})\right].$ (25)
Reference [33] further proved the following inequality:
$\mathbb{E}_{\xi}\left[\sup_{{h}_{j}\in\mathcal{H}_{j}}\sum_{m\in\mathcal{M}}\sum_{j=1}^{{N_{\mathrm{B}}}}\xi^{(m,j)}{h}_{j}(\mathbf{x}^{(m)})\right]$
$\leq\sum_{j=1}^{{N_{\mathrm{B}}}}\mathbb{E}_{\xi}\left[\sup_{{h}_{j}\in\mathcal{H}_{j}}\sum_{m\in\mathcal{M}}\xi^{(m,j)}{h}_{j}(\mathbf{x}^{(m)})\right].$ (26)
Then, this empirical Rademancher complexity is bounded by:
^$\widehat{\mathcal{R}}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})$ $\leq\sqrt{2}C_{\ell}\sum_{j=1}^{{N_{\mathrm{B}}}}\mathbb{E}_{\xi}\left[\sup_{{h}_{j}\in\mathcal{H}_{j}}\sum_{m\in\mathcal{M}}\xi^{(m,j)}{h}_{j}(\mathbf{x}^{(m)})\right],$ (27)
where $\mathcal{X}=\{\mathbf{x}_{i},\forall m\in\mathcal{M}\}$. Observing that the empirical Rademancher complexity of the hypothesis class $\mathcal{H}_{j}$ on dataset $\mathcal{X}$ is defined as:
^$\widehat{\mathcal{R}}_{\mathcal{X}}(\mathcal{H}_{j})=\mathbb{E}_{\xi}\left[\sup_{{h}_{j}\in\mathcal{H}_{j}}\sum_{m\in\mathcal{M}}\xi^{(m,j)}{h}_{j}(\mathbf{x}^{(m)})\right],$ (28)
Eq. (27) can be further expressed as:
^$\widehat{\mathcal{R}}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})\leq\sqrt{2}C_{\ell}\sum_{j=1}^{{N_{\mathrm{B}}}}\widehat{\mathcal{R}}_{\mathcal{X}}(\mathcal{H}_{j}).$ (29)
Before calculating ^$\widehat{\mathcal{R}}_{\mathcal{X}}(\mathcal{H}_{j})$, we first introduce the positive homogeneity of ReLU $\phi(\cdot)$ as follows:
$a\cdot\phi(\mathbf{x})=\phi(a\cdot\mathbf{x}),\ \forall a\geq 0.$ (30)
Meanwhile, from the L2 regularization, we have:
$\|\mathbf{w}^{k}_{j}\|_{2}\leq W^{\mathrm{max}}\quad\Rightarrow\quad\|\mathbf{w}^{k}_{j}\|_{\infty}\leq\sqrt{W^{\mathrm{max}}}.$ (31)
We further define a new variable $\overline{\mathbf{w}}^{k}_{j}=\frac{\mathbf{w}^{k}_{j}}{{\|\mathbf{w}^{k}_{j}\|_{2}}}$. Obviously, we have $\|\overline{\mathbf{w}}^{k}_{j}\|_{2}=1$. Then, according to the definition of $h_{j}(\mathbf{x})$, the empirical Rademancher complexity ^$\widehat{\mathcal{R}}_{\mathcal{X}}(\mathcal{H}_{j})$ can be expressed as:
^$\widehat{\mathcal{R}}_{\mathcal{X}}(\mathcal{H}_{j})=\frac{1}{M}\mathbb{E}_{\xi}\left[\sup_{\mathbf{h}^{K-2}\in\mathcal{H}^{K-2}}\sum_{m\in\mathcal{M}}\xi^{(m)}\left(\sum_{j^{\prime}\in\mathcal{J}^{K-1}}w^{K}_{j,j^{\prime}}\right.\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.\left.\cdot\phi\left((\mathbf{w}^{K-1}_{j^{\prime}})^{\intercal}\phi(\mathbf{h}^{K-2}(\mathbf{x}^{(m)}))\right)\right)\right]$
$=\frac{1}{M}\mathbb{E}_{\xi}\left[\sup_{\mathbf{h}^{K-2}\in\mathcal{H}^{K-2}}\sum_{j^{\prime}\in\mathcal{J}^{K-1}}\underbrace{w^{K}_{j,j^{\prime}}\left\|\mathbf{w}^{K-1}_{j^{\prime}}\right\|_{2}}_{a_{j^{\prime}}}\right.$
$\quad\quad\quad\quad\quad\quad\left.\cdot\underbrace{\sum_{m\in\mathcal{M}}\xi^{(m)}\phi\left((\overline{\mathbf{w}}^{K-1}_{j^{\prime}})^{\intercal}\phi(\mathbf{h}^{(K-2)})\right)}_{b_{j^{\prime}}}\right],$ (32)
where $w^{K}_{j,j^{\prime}}$ is the $j^{\prime}$ component of vector $\mathbf{w}^{K}_{j}$. The second inequality holds due to the homogeneity of ReLU. Since $\sum_{j^{\prime}}a_{j^{\prime}}b_{j^{\prime}}\leq\sum_{j^{\prime}}|a_{j^{\prime}}|\max_{j^{\prime\prime}}|{b_{{j^{\prime\prime}}}}|$ holds for any $a_{j^{\prime}}$ and $b_{j^{\prime}}$, the value of ^$\widehat{\mathcal{R}}_{\mathcal{X}}(\mathcal{H}_{j})$ is bounded by:
^$\widehat{\mathcal{R}}_{\mathcal{X}}(\mathcal{H}_{j})\leq\frac{1}{M}\mathbb{E}_{\xi}\left[\sup_{\mathbf{h}^{K-2}\in\mathcal{H}^{K-2}}\sum_{j^{\prime}\in\mathcal{J}^{K-1}}\underbrace{\left|w^{K}_{j,j^{\prime}}\right|\left\|\mathbf{w}^{K-1}_{j^{\prime}}\right\|_{2}}_{|a_{j^{\prime}}|}\right.$
$\quad\quad\quad\quad\left.\cdot\underbrace{\max_{j^{\prime\prime}}\left|\sum_{m\in\mathcal{M}}\xi^{(m)}\phi\left((\overline{\mathbf{w}}^{K-1}_{j^{\prime\prime}})^{\intercal}\phi(\mathbf{h}^{K-2})\right)\right|}_{\max_{j^{\prime\prime}}|{b_{j^{\prime\prime}}}|}\right].$ (33)
According to (31), we further have:
^$\widehat{\mathcal{R}}_{\mathcal{X}}(\mathcal{H}_{j})$
$\leq\frac{(W^{\mathrm{max}})^{\frac{3}{2}}}{M}\mathbb{E}_{\xi}\left[\sup_{\begin{subarray}{c}\mathbf{h}^{K-2}\in\mathcal{H}^{K-2},\\ \left\|\overline{\mathbf{w}}\right\|_{2}=1,\end{subarray}}\left|\sum_{m\in\mathcal{M}}\xi^{(m)}\phi\left(\overline{\mathbf{w}}^{\intercal}\phi(\mathbf{h}^{K-2})\right)\right|\right]$ (36)
$\leq\frac{(W^{\mathrm{max}})^{\frac{3}{2}}}{M}\mathbb{E}_{\xi}\left[\sup_{\begin{subarray}{c}\mathbf{h}^{K-2}\in\mathcal{H}^{K-2},\\ \left\|\overline{\mathbf{w}}\right\|_{2}\leq 1,\end{subarray}}\left|\sum_{m\in\mathcal{M}}\xi^{(m)}\phi\left(\overline{\mathbf{w}}^{\intercal}\phi(\mathbf{h}^{K-2})\right)\right|\right],$ (39)
In the first inequality above, we replace $\overline{\mathbf{w}}^{K-1}_{j^{\prime}}$ by a new variable $\overline{\mathbf{w}}$, which has the same dimension as $\mathbf{w}^{K-1}_{j^{\prime}}$. this replacement is equivalent because $\overline{\mathbf{w}}^{K-1}_{j^{\prime}}$ and $\overline{\mathbf{w}}$ are only constrained by $\left\|\overline{\mathbf{w}}\right\|_{2}=1$ and $\|\mathbf{w}^{K-1}_{j^{\prime}}\|_{2}=1$, respectively. The second inequality holds because it enlarges the hypothesis class. Based on [Lemma 5.12] in [34], we can remove the absolute operator in (39) and obtain:
^$\widehat{\mathcal{R}}_{\mathcal{X}}(\mathcal{H}_{j})$
$\leq\frac{2(W^{\mathrm{max}})^{\frac{3}{2}}}{M}\mathbb{E}_{\xi}\left[\sup_{\begin{subarray}{c}\mathbf{h}^{K-2}\in\mathcal{H}^{K-2},\\ \left\|\overline{\mathbf{w}}\right\|_{2}\leq 1,\end{subarray}}\sum_{m\in\mathcal{M}}\xi^{(m)}\phi\left(\overline{\mathbf{w}}^{\intercal}\phi(\mathbf{h}^{K-2})\right)\right]$ (42)
$\leq\frac{2(W^{\mathrm{max}})^{\frac{3}{2}}}{M}\mathbb{E}_{\xi}\left[\sup_{\begin{subarray}{c}\mathbf{h}^{K-2}\in\mathcal{H}^{K-2},\\ \left\|\overline{\mathbf{w}}\right\|_{2}\leq 1,\end{subarray}}\sum_{m\in\mathcal{M}}\xi^{(m)}\left(\overline{\mathbf{w}}^{\intercal}\phi(\mathbf{h}^{K-2})\right)\right]$ (45)
$=2(W^{\mathrm{max}})^{\frac{3}{2}}\widehat{\mathcal{R}}_{\mathcal{X}}(\overline{\mathcal{H}}^{K-1}),$ (46)
where the definition of class ¯$\overline{\mathcal{H}}^{k}$ is almost the same as ${\mathcal{H}}^{k}$ defined in (22) except for replacing the L2 regulation constraint $\|\mathbf{w}^{(k)}_{j}\|_{2}\leq W^{\mathrm{max}}$ with $\left\|\overline{\mathbf{w}}\right\|_{2}\leq 1$. The second inequality holds because $\phi(\cdot)$ is 1-Lipschitz continuous. The last equality holds due to the definition of the empirical Rademancher complexity. In a similar way, we can also obtain:
^$\widehat{\mathcal{R}}_{\mathcal{X}}(\overline{\mathcal{H}}^{k})\leq 2W^{\mathrm{max}}\widehat{\mathcal{R}}_{\mathcal{X}}(\overline{\mathcal{H}}^{k-1}),\forall k\in\{2,\cdots,K-1\},$ (47)
where the coefficient is $2W^{\mathrm{max}}$ instead of $2(W^{\mathrm{max}})^{\frac{3}{2}}$ because the infinity norm of $\overline{\mathbf{w}}$ is bounded by one instead of $\sqrt{W^{\mathrm{max}}}$. By substituting (46) and (47) recursively, we obtain:
^$\widehat{\mathcal{R}}_{\mathcal{X}}(\mathcal{H}_{j})$ $\leq\sqrt{W^{\mathrm{max}}}\cdot 2W^{\mathrm{max}}\cdot\widehat{\mathcal{R}}_{\mathcal{X}}(\overline{\mathcal{H}}^{K-1})$
$\leq\sqrt{W^{\mathrm{max}}}\cdot(2W^{\mathrm{max}})^{2}\cdot\widehat{\mathcal{R}}_{\mathcal{X}}(\overline{\mathcal{H}}^{K-2})$
$\cdots$
$\leq\sqrt{W^{\mathrm{max}}}\cdot(2W^{\mathrm{max}})^{K-1}\cdot\widehat{\mathcal{R}}_{\mathcal{X}}(\overline{\mathcal{H}}^{1}).$ (48)
The Rademancher complexity of the loss is the expectation of the empirical one over the probability distribution of $\mathbf{x}$. By substituting (48) into (29), we can bound this Rademancher complexity via:
$\mathcal{R}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})$ $=\mathbb{E}_{\mathbf{x}}\left(\widehat{\mathcal{R}}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})\right)$
$\leq\sqrt{2}C_{\ell}\cdot\mathbb{E}_{\mathbf{x}}\left(\widehat{\mathcal{R}}_{\mathcal{X}}(\mathcal{H}_{j}^{K})\right)$
$\leq\sqrt{2}C_{\ell}\cdot\sqrt{W^{\mathrm{max}}}\cdot(2W^{\mathrm{max}})^{K-1}\mathbb{E}_{\mathbf{x}}\left(\widehat{\mathcal{R}}_{\mathcal{X}}(\overline{\mathcal{H}}^{1})\right)$
$\leq\sqrt{2}C_{\ell}\cdot\sqrt{W^{\mathrm{max}}}\cdot(2W^{\mathrm{max}})^{K-1}\mathcal{R}_{\mathcal{S}}(\overline{\mathcal{H}}^{1}),$ (49)
where ¯$\overline{\mathcal{H}}^{1}=\{\mathbf{x}\rightarrow\overline{\mathbf{w}}^{\intercal}\mathbf{x},\|\overline{\mathbf{w}}\|_{2}\leq 1\}$. According to [Theorem 3] in [35], its Rademancher complexity is bounded by:
$\mathcal{R}_{\mathcal{S}}(\overline{\mathcal{H}}^{1})\leq\frac{X^{\mathrm{max}}}{\sqrt{M}}.$ (50)
By substituting (50) into (49), we obtain an upper bound of the Rademancher complexity $\mathcal{R}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})$, as follows:
$\mathcal{R}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})\leq\frac{(2W^{\mathrm{max}})^{K-\frac{1}{2}}C_{\ell}X^{\mathrm{max}}}{\sqrt{M}}.$ (51)
B-3 PAC upper bound of the excess cost
The following Lemma gives a PAC bound for the excess cost of the hypothesis $\widehat{\mathbf{h}}$ based on uniform convergence.
Lemma 2 (Generalization bound via uniform convergence [36]).
When Assumptions A1 and A2 hold, with probability at least $1-\delta$ for any small $\delta\in(0,1)$, the excess cost of hypothesis $\widehat{\mathbf{h}}$ is bounded by:
$\Delta L(\widehat{\mathbf{h}})\leq 4\mathcal{R}_{\mathcal{S}}(\mathcal{L}_{\mathbf{h}})+\sqrt{\frac{2\ln(2/\delta)}{M}}.$ (52)
By substituting (51) into (52), Theorem 1 can be proved.